# NAMO-LLM: Efficient Navigation Among Movable Obstacles with Large Language Model Guidance

### v1

@@ IV-A Setting Up Comparative Experiments
IV-B Comparative Experiments against NAMO Baselines
In what follows, we present six case studies that differ in their complexity in terms of the total number $n$ of movable obstacles, the initial configuration of the obstacles, and the minimum horizon, denoted by $H_{\text{min}}$, of a plan $\tau$ required to reach ${\mathcal{P}}_{g}$. In our method, NAMO-LLM$(p_{\text{rand}},p_{\text{obs}})$, we employ distinct parameter settings for $p_{\text{rand}}$ and $p_{\text{obs}}$. The following results are obtained using the pre-trained GPT 4o model.
@@ IV-B Comparative Experiments against NAMO Baselines
[…]
@@
[…]
@@
Case Study I ($n=50,H_{\text{min}}=1$): In this case study, we consider a $50\times 50$ environment with $n=50$ movable obstacles. The initial configurations of the obstacles and the robot, as well as the desired robot configuration, are depicted in Fig. 5. The environment is designed such that a plan with horizon $H_{\text{min}}=1$ exists to reach ${\mathcal{P}}_{g}$, meaning that the minimum number of obstacles that must be reconfigured to reach ${\mathcal{P}}_{g}$ is $1$. We test NAMO-LLM for various values of the parameters $p_{\text{rand}}$, $p_{\text{obs}}$, and $K$. The comparative results are summarized in Table I. Observe that our method significantly outperforms both baselines in terms of both the runtime required to compute the first feasible plan and the horizon $H$ of that plan. For instance, the runtimes of the search- and sampling-based methods are $99.05$ and $14.94$ seconds, respectively, while our method can compute the first feasible plan significantly faster; e.g., the average runtime for $K=1$, $p_{\text{rand}}=0.8$, and $p_{\text{obs}}=0.8$ is $2.88$ seconds. The average horizon of the plan computed by our method is consistently less than $4$, which is comparable to that of the sampling-based baseline, while the horizon of the search-based one is $27$. We note that the performance of our method does not change significantly with respect to $K\in\{1,2,3\}$. We attribute this to the fact that there are multiple candidate obstacles near ${\mathcal{P}}_{g}$ that can be removed to get access to it and that the employed LLM is capable of identifying them.
Case Study II ($n=100,H_{\text{min}}=1$): In this case study, we consider a $65\times 65$ environment containing $n=100$ movable obstacles, where ${\mathcal{P}}_{g}$ can be reached by removing at least one obstacle; see Fig 5. The comparative results are shown in Table II. Due to space limitations, we present results for our method only with $K=1$. Observe that the performance gap between our method and the baselines increases as $n$ grows. For instance, when configured with $K=1$, $p_{\text{rand}}=0.8$, and $p_{\text{obs}}=0.8$, our method finds the first feasible plan in $7.80$ seconds with an average horizon of $H=3.04$. In contrast, [22] takes $42.78$ seconds to compute a plan with a slightly longer horizon ($H=4.57$), while [18] requires $876.36$ seconds and produces a plan with $H=65$.
@@
Case Study III ($n=50,H_{\text{min}}=2$): Here we consider a $50\times 50$ environment with $n=50$ movable obstacles, where reaching ${\mathcal{P}}_{g}$ requires removing at least two obstacles; see Fig 5. The results in Table II show that the performance gap between our method and the baselines further increases as the minimum required horizon $H_{\text{min}}$ grows.
Case Study IV ($n=100,H_{\text{min}}=2$): This case study examines a more complex environment with dimensions $65\times 65$ and $n=100$ movable obstacles, where reaching ${\mathcal{P}}_{g}$ requires removing at least two obstacles; see Fig 5. The results are summarized in Table II, showing that the performance gap further increases as both $n$ and the minimum horizon $H_{\text{min}}$ increase. For instance, our method, when configured with $K=1$, $p_{\text{rand}}=0.8$, and $p_{\text{obs}}=0.8$, finds the first feasible plan in $17.12$ seconds with an average horizon of $H=4.5$. In contrast, the sampling-based approach takes $408.68$ seconds to compute a plan with a longer horizon of $7.41$, while the search-based method requires $877.97$ seconds and produces a plan with $H=65$.
[…]
@@
Case Study V ($n=50,H_{\text{min}}=6$): In this case study, we consider a $40\times 40$ environment with $n=50$ movable obstacles, where reaching ${\mathcal{P}}_{g}$ requires removing at least six obstacles (see Fig. 6). This scenario is significantly more complex due to the larger number of obstacles that need to be relocated. This complexity is also reflected in the increased runtimes reported in Table III, compared to the previous case studies. Our method, when configured with $K=1$, $p_{\text{rand}}=0.8$, and $p_{\text{obs}}=0.8$, computed the first feasible plan in $48.22$ seconds, with an average horizon of $H=26.48$. In contrast, [22] failed to find a feasible plan within $6$ hours while [18] can identify a feasible plan within $141.91$ seconds with $H=42$.
@@
Case Study VI ($n=10,H_{\text{min}}=3$): Finally, we consider a more complex scenario to illustrate the utility of parameter $K$, where—unlike previous case studies—some manipulable obstacles are blocked by others and cannot be relocated until the blocking ones are moved. Specifically, we consider a $20\times 20$ environment with $n=10$ movable obstacles, where reaching ${\mathcal{P}}_{g}$ requires relocating the manipulable obstacle $o_{1}$. However, unlike Case Studies I-IV, there do not exist collision-free paths to any possible configuration $c_{1}^{\prime}$ unless both $o_{8}$ and $o_{2}$ are moved first; see Fig 6. Thus, in this case, at least $3$ obstacles must be relocated to reach ${\mathcal{P}}_{g}$. The comparative results are shown in Table IV. Unlike Case Study I, the performance of our method improves significantly for $K\in\{2,3\}$ compared to $K=1$. This occurs because, with $K=1$, the LLM often prioritizes removing $o_{1}$ first (i.e., $o_{\text{LLM}}=o_{1}$ in (2)), even though it cannot be moved due to being blocked by $o_{8}$ and $o_{2}$. Consequently, any configuration $c_{1}^{\prime}$ sampled for $o_{1}$ is rejected by line 2 in Alg. 2, slowing down the overall algorithm. However, when $K>1$, the LLM generates a set of obstacles that need to be removed to access ${\mathcal{P}}_{g}$, which may include $o_{2}$ and/or $o_{8}$. Randomly selecting an obstacle from this set helps the algorithm avoid getting stuck due to persistent incorrect LLM recommendations and, consequently, improves performance. Also, observe that our method outperforms [22], which relies on random exploration, in terms of runtime—even when configured with $K=1$. We attribute this to the fact that the LLM occasionally provides correct guidance (i.e., recommending the removal of $o_{2}$ and $o_{8}$ before $o_{1}$) after some redundant actions have been applied (e.g., relocating $o_{3}$ first), as these actions influence the input prompt to the LLM. Overall, our method outperforms both baselines when $K$ is $2$ or $3$. Observe in Table IV that [18] eventually moved all obstacles to the storage area leading to plans with $H=10$.
@@
Summary of Comparisons: Overall, the performance gap between NAMO-LLM and the baselines increases as $n$ or $H_{\text{min}}$ grows. In general, we observed that our method tends to achieve its lowest runtime for computing the first feasible plan when configured with $p_{\text{rand}}=p_{\text{obs}}=0.8$ and $K=1$; if manipulable obstacles are not blocked by other obstacles; see e.g., Case Studies I–V. We attribute this to the LLM’s ability to correctly reason about which obstacle to remove in these scenarios. However, when $p_{\text{rand}}=0.2$ and $p_{\text{obs}}=0.8$, the generated plans tend to have a shorter horizon $H$. This occurs because a smaller $p_{\text{rand}}$ enables the algorithm to select tree nodes for expansion more uniformly, leading to better-quality first plans, albeit with slightly higher runtimes. We also observed that pre-trained LLMs may occasionally make infeasible recommendations regarding which manipulable obstacle to relocate, especially when these obstacles are blocked by others and cannot be repositioned until the blocking ones are moved first; see e.g., Case Study VI. In such cases, setting $K>1$ is beneficial, as it prevents the algorithm from getting stuck due to poor LLM recommendations. Finally, we note that these results depend on the selected LLM.
@@
[…]
@@
[…]
@@
[…]
@@
TABLE V: Effect of LLM Model on average tree size, runtimes, and horizon. The line ‘GPT 4o’ reports the best results of our method when paired with GPT 4o.
IV-C Comparative Results with an Alternative LLM
In this section, we demonstrate the performance of our method when paired with the Gemini-1.5-flash model on Case Studies IV ($n=100$, $H_{\text{min}}=2$) and V ($n=50$, $H_{\text{min}}=6$). The comparative results are shown in Table V. In Case Study IV, our method outperforms both baselines in terms of runtime but produces longer plans than the sampling-based baseline (while still shorter than those of the search-based one). In Case Study V, however, our approach achieves better performance than both baselines in terms of both runtime and plan horizon $H$. Overall, we observe that switching from GPT-4o to Gemini-1.5-flash tends to improve the performance of our algorithm in scenarios where manipulable obstacles are blocked by other obstacles (see Case Study V) and deteriorate its performance in cases with large number of movable obstacles (see Case Study IV). Nevertheless, our method still remains more scalable than the baselines.
@@ V Conclusion
In this paper, we proposed NAMO-LLM, a probabilistically complete sampling-based planner for NAMO problems. Through comparative experiments, we demonstrated its scalability and efficiency in highly cluttered environments. Future work will focus on extending NAMO-LLM to settings with obstacles of unknown location and movability.

### v2

@@ IV-B Comparisons with Search- and Sampling-Based Baselines
TABLE I: Case Study I: Average tree size $|{\mathcal{V}}|$, runtimes, and horizon $H$.
@@
TABLE II: Case Study II-IV: Average tree size, runtimes, and horizon.
@@
Case Study I ($n=50,H_{\text{min}}=1$): We consider a $50\times 50$ environment with $n=50$ movable obstacles admitting a feasible plan with $H_{\text{min}}=1$; see Fig. 5. We compared NAMO-LLM for various values of $p_{\text{rand}}$, $p_{\text{obs}}$, and $K$ against the baselines; see Table I. Observe that our method significantly outperforms both baselines in terms of both the runtime required to compute the first feasible plan and the horizon $H$ of that plan. For instance, NAMO-LLM tends to be 30x and 5x faster than [16] and [12], respectively. Observe that the performance of our method does not change significantly with respect to $K\in\{1,2,3\}$. We attribute this to the fact that there are multiple candidate obstacles near the goal area ${\mathcal{P}}_{g}$ that can be removed to get access to it and that the employed LLM is capable of identifying them.
In Case Studies II-V, we present results for our method with $K=1$ combined with $p_{\text{rand}}=p_{\text{obs}}=0.8$, and $p_{\text{rand}}=0.2$, $p_{\text{obs}}=0.8$.
@@
In Case Study II ($n=100,H_{\text{min}}=1$), Case Study III ($n=50,H_{\text{min}}=2$) and Case Study IV ($n=100,H_{\text{min}}=2$), we similarly observe that the performance gap between our method and the baselines further increases as the minimum required horizon $H_{\text{min}}$ grows and/or $n$ increase. The environments associated with these case studies can be found in Fig 5 while the results are shown in Table II.
@@
Case Study V ($n=50,H_{\text{min}}=6$): We consider a $40\times 40$ environment; see Fig. 6. This scenario is significantly more complex due to the larger number of obstacles that need to be relocated; this is also reflected in the increased runtimes reported in Table III, compared to the previous case studies. Our method, when configured with $K=1$, $p_{\text{rand}}=p_{\text{obs}}=0.8$ computed the first feasible plan in $48.22$ secs, with an average horizon of $H=26.48$. In contrast, [16] failed to find a feasible plan within $6$ hours while [12] can identify a feasible plan within $141.91$ secs with $H=42$.
Figure 6: Case Study V & VI: Initial configurations of the obstacles (red polygons). The green area corresponds to ${\mathcal{P}}_{g}$.
@@
Case Study VI ($n=10,H_{\text{min}}=3$): Finally, we consider a more complex scenario to illustrate the utility of parameter $K$. Unlike previous case studies, here some manipulable obstacles are blocked by others. Specifically, we consider a $20\times 20$ environment with $n=10$ movable obstacles, where reaching $\mathcal{P}_{g}$ requires relocating $o_{1}$, but no collision-free path to any ${\mathbf{c}}_{1}^{\prime}$ exists until $o_{8}$ and $o_{2}$ are moved; see Fig. 6. Table IV shows that, unlike Case I, performance improves significantly for $K\in\{2,3\}$ compared to $K=1$. This occurs because because with $K=1$ the LLM often prioritizes $o_{1}$, which is blocked by $o_{8}$ and $o_{2}$; any ${\mathbf{c}}_{1}^{\prime}$ sampled for $o_{1}$ is rejected at line 2 in Alg. 2, slowing progress. However, for $K>1$, the LLM outputs a set of candidate obstacles (often including $o_{2}$ and/or $o_{8}$), and random selection from this set helps avoid persistent incorrect recommendations. Even with $K=1$, our method outperforms [16] in runtime because the LLM occasionally recommends $o_{2}$ and $o_{8}$ after some exploratory moves (e.g., relocating $o_{3}$), which alters its prompt. Overall, our method surpasses both baselines for $K=2,3$; note in Table IV that [12] moved all obstacles to storage since $H=10$.
@@
Summary: The performance gap between NAMO-LLM and the baselines increases as $n$ or $H_{\text{min}}$ grows. Our method achieves its lowest runtime for computing the first feasible plan when $p_{\text{rand}}=p_{\text{obs}}=0.8$ and $K=1$, provided manipulable obstacles are not blocked (Cases I–V). This is due to the LLM’s ability to identify the correct obstacle to remove in these scenarios. When $p_{\text{rand}}=0.2$ and $p_{\text{obs}}=0.8$, plans typically have shorter horizons, since more uniform node expansion leads to higher-quality first plans, albeit with slightly longer runtimes. We also observed that LLMs may produce infeasible recommendations when manipulable obstacles are blocked (e.g., Case VI). In such cases, setting $K>1$ prevents the algorithm from stalling on poor LLM predictions. Finally, across all case studies, $2.77\%$ of LLM outputs are invalid (format or content errors) on average, while $\sim 70\%$ of which are due to content errors; these are handled via uniform sampling as discussed in Section III-C.
@@
TABLE III: Case Study V: Average tree size, runtimes, and horizon.
@@
TABLE IV: Case Study VI: Average tree size, runtimes, and horizon.
@@
MiniGPT4o(0.2, 0.8) 19.98 81.45 5.62
MiniGPT4o(0.8, 0.8) 10.18 59.72 6.96
Gemini(0.2, 0.8) 84.35 233.94 15.8
Gemini(0.8, 0.8) 24.75 104.30 17.2
Case V NAMO-SA [12] N/A 141.91 42
@@
MiniGPT4o(0.2, 0.8) 52.1 67.61 13.23
MiniGPT4o(0.8, 0.8) 20.3 27.13 13.57
Gemini(0.2, 0.8) 27.70 21.03 9.6
Gemini(0.8, 0.8) 15.15 14.24 12.15
TABLE V: Performance comparison of different models in Cases IV and V, showing average tree size, runtimes, and horizon. The line ‘GPT 4o’ reports the best results of our method when paired with GPT 4o. Lines of the form ‘LLM($p_{\text{rand}},p_{\text{obs}}$)’ refer to our method when coupled with the respective LLM.
IV-C Comparative Results with Alternative LLMs & VLMs
We evaluate NAMO-LLM with GPT-4o, GPT-4o-mini, and Gemini-1.5-Flash on Cases IV and V; see the Table V). Switching among these models affects performance, but our method remains more scalable than the baselines. In Case Study IV, it outperforms both baselines in runtime for all tested LLMs, though with the Gemini model it produces longer plans than [16] (while still shorter than those of [12]). In Case Study V, our method outperforms both baselines in both metrics for every LLM tested. We also observed that the performance of our method deteriorates when it is paired with VLMs (GPT 4o). For instance, our average runtime and horizon on Case IV with $p_{\text{rand}}=p_{\text{obs}}=0.8$ is $252.21$ secs and $32.8$, respectively, while it fails to address Case V within $1,000$ secs. The VLM is prompted with an overhead image of the environment, replacing the textual description of the environment. We attribute this result to potentially limited VLM capabilities in object recognition and counting tasks.
IV-D Comparisons with LLM-based Planners
We compare NAMO-LLM against the LLM-based planner from Section IV-A, using GPT-4o. The baseline’s success rates in Cases I–VI are $86\%$, $92\%$, $78\%$, $90\%$, $54\%$, and $0\%$, respectively, while NAMO-LLM achieves $100\%$ in all. GPT-4o alone is competitive in simpler tasks (Cases I–IV) but drops sharply in complex ones (Cases V–VI). [^2] Other LLMs showed the same trend; e.g., GPT-4o-mini achieved 81%, 70%, 92%, 82%, 55%, and 0% for Cases I–VI, respectively. For Cases I–IV, both methods produce plans with comparable horizons. As complexity increases (Cases V–VI), NAMO-LLM outperforms the baseline. Across Cases I–VI, the best (average) horizons of NAMO-LLM are $2.5$, $2.75$, $3.70$, $4.16$, $19.65$, and $8.11$, compared to $2.68$, $2.7$, $3.4$, $3.75$, $52.88$, and N/A for the baseline. For runtimes, the baseline is slightly faster in simpler tasks but fails on more complex ones (Case V): NAMO-LLM achieves $2.88$, $7.80$, $5.91$, $17.12$, and $31.05$ secs across Cases I–V, compared to $2.73$, $6.47$, $5.37$, $13.17$, and N/A for the baseline. Overall, these comparisons show that our algorithm consistently outperforms the baseline in terms of task success rates, particularly in complex scenarios (e.g., Case V). The main advantage of the baseline is that, if it produces a plan, the plan may be found slightly faster especially in simpler tasks (Cases I–V).
IV-E Hardware Validation Example
To demonstrate the practical feasibility of transferring NAMO-LLM plans from simulation to real-world execution, we validated the method on a TurtleBot3 Waffle Pi with a 4-DOF OpenManipulator-X, operating in a 3D Mini-City platform; see Fig. 1 and the supplemental material. The robot performed state estimation, navigation, and manipulation using off-the-shelf ROS packages. The experiment involved $n=19$ movable obstacles (traffic cones, construction blocks) and several fixed objects (buildings, pedestrian islands), with a minimal horizon $H_{\min}=2$. To facilitate computation of free-space components ${\mathcal{N}}(t)$ we over-approximated the robot as a disk of radius $r=15cm$ and the obstacles as polygons. Our algorithm found the first feasible path in $1.14$ secs on average and the optimal $H_{\min}$ plan in $7.32$ secs.
@@ V Conclusion
This paper proposed a probabilistically complete planner for NAMO problems. Comparative experiments verified its efficiency in cluttered environments. Future work will focus on extending to settings with unknown obstacle location and movability, and developing variants that avoid explicit C-space computation to better accommodate complex geometries and high-dimensional configuration spaces.